# Unraveling Indirect In-Context Learning Using Influence Functions

### v1

@@ 4.5 Computational Complexity
We present the worst case time complexity for inference of ours and related method in Table 4. As can be observed, our methods are comparable, if not more efficient than the other baselines. Note that the Sur methods require an additional fine-tuning step on a smaller surrogate model before the gradients are extracted, which the Pre methods do not. Furthermore, note that TracIn as an influence method is much faster than Hessian-based approaches (e.g. DataInf) as it assumes that the Hessian is the identity matrix. While this leads to more efficient influence computation, it comes at the cost of lower estimation performance, as our results with TracIn also show.
@@ 5 Related Work
IFs have been utilized with great success in a number of application scenarios (e.g. classification Chhabra et al. (2024a); Koh and Liang (2017), generative models Kwon et al. (2023); Schioppa et al. (2022); Grosse et al. (2023), active learning Chhabra et al. (2024b); Liu et al. (2021b), etc.). Moreover, while some recent works have considered using influence for selecting direct demonstrations (Nguyen and Wong, 2023; Van et al., 2024), neither of them has consider their effect on inductive bias selection in the indirect ICL setting, which is the focus of our work.
@@ 6 Conclusion
We formalize a new paradigm for generalized In-Context Learning, which we call Indirect In-Context Learning. We analyze two different real-world Indirect ICL settings and propose effective demonstration selection strategies for these scenarios. We explore using Influence Functions (IFs) to leverage the informativeness of the samples in the demonstration pool and the models’ task inductive bias. We find that combining a surrogate model-based IF approach with BertScore performs better when there are an overwhelming majority of irrelevant tasks in the candidate pool. We also find that reweighting the surrogate model-based IF scores with traditional metric scores can be helpful in the case where noisy demonstrations are present. Future work will aim to augment the Pretrained Gradient approach by using better/larger LLMs or finetuning the LLMs.
@@
Appendix A Indirect ICL Results for LLM based Influence
A.1 Full Results for Mixture of Tasks
@@
Following are the full results for the Mixture of Tasks setting. For $k=3$ shots in Tables 5, 6, and 7. For $k=5$ shots in Table 8.
@@
Table 5: Performance across different datasets and demonstration selection methods ($k=3$ shots). The datasets are sampled from sub-tasks of the MMLU and BigBench datasets for Llama-2-13b-chat.
@@
Dataset Rand BSR Cos BM25 PreD SurD Sur[D,BSR] Sur[D,Cos] medical-genetics 80.60 87.00 84.00 79.00 76.00 81.00 86.00 82.00 prof-psychology 68.30 70.00 73.00 65.50 65.00 68.50 72.00 68.50 formal-logic 60.16 59.52 60.32 58.73 61.11 59.52 57.14 55.56 moral-disputes 81.00 78.00 79.50 80.50 78.50 76.00 80.50 76.50 public-relations 72.18 79.09 80.91 73.64 71.82 75.45 79.09 79.09 comp-security 76.80 76.00 80.00 76.00 76.00 80.00 76.00 76.00 astronomy 80.26 80.26 78.95 80.92 74.34 79.61 80.26 78.95 abstract-algebra 57.00 58.00 62.00 55.00 57.00 47.00 72.00 72.00 nutrition 75.50 77.50 79.00 78.00 77.50 77.00 79.50 81.50 high-school-biology 76.70 76.50 78.00 76.50 73.50 79.00 80.50 76.50 formal-fallacies 47.25 52.00 50.00 49.50 47.00 56.50 47.50 50.00 tracking-3 40.20 44.00 39.00 45.00 40.00 37.00 43.50 39.00 Average 68.00 69.87 70.39 68.19 66.48 68.04 71.16 69.63
@@
Table 6: Performance across different datasets and demonstration selection methods ($k=3$ shots). The datasets are sampled from sub-tasks of the MMLU and BigBench datasets for Mistral-7b-v3
@@
Table 7: Performance across different datasets and demonstration selection methods ($k=3$ shots). The datasets are sampled from sub-tasks of the MMLU and BigBench datasets for Zephyr-7b-beta
@@
Table 8: Performance across different datasets and demonstration selection methods ($k=5$ shots)
@@
A.2 TracIn Results
Here we provide results for the TracIn method of Influence Computation for $k=3$ shots in Tables 9, 10, and 11. We also provide results for $k=5$ shots in Table 12.
Table 9: Performance across different datasets and different TracIn Influence methods for $k=3$ shots for Llama2-13b-chat
@@ A.2 TracIn Results
Table 10: Performance across different datasets and different TracIn Influence methods for $k=3$ shots for Mistral-7b-v0.3
@@
Table 11: Performance across different datasets and different TracIn Influence methods for $k=3$ shots for Zephyr-7b-beta
@@
Table 12: Performance across different datasets and different TracIn Influence methods for $k=5$ shots for Llama2-13b-chat.
@@
A.3 Pretrained Gradient Results
Here we provide results for pretrained gradients method of computing IF. These can be found in for $k=3$ shots in Tables 13, 14 and 15 and for $k=5$ shots in Table 16.
@@ A.3 Pretrained Gradient Results
Table 14: Performance across different datasets and Pre-training based demonstration selection methods ($k=3$ shots) for Mistral-7b-v0.3.
@@
Table 15: Performance across different datasets and Pre-training based demonstration selection methods ($k=3$ shots) for Zephyr-7b-beta.
@@
Table 16: Performance across different datasets and Pre-training based demonstration selection methods ($k=5$ shots) for Llama2-13b-chat.
@@
Appendix B Noisy ICL Results for different parameters
@@
Here we provide the results for different noise levels in the dataset and the different hyperparameters we tested in Table 17.

### v2

@@ 4.5 Computational Complexity
We present the worst case time complexity (for inference) for our methods and related baselines in Table 2. As can be observed, our methods are comparable, if not more efficient than the other baselines. Note that the Sur methods require an additional fine-tuning step on a smaller surrogate model before the gradients are extracted, which the Pre methods do not. Furthermore, note that TracIn as an influence method is much faster than Hessian-based approaches (e.g. DataInf) as it assumes that the Hessian is the identity matrix. While this leads to more efficient influence computation, it comes at the cost of lower estimation performance, as our results with TracIn also show. Additionally, we present the maximum GPU memory consumption while performing demonstration selection in Appendix E, scalability to large models in Appendix F and scalability to large datasets in Appendix G
@@ 5 Related Work
IFs have been utilized with great success in a number of application scenarios (e.g. classification Chhabra et al. (2024a); Koh and Liang (2017), generative models Kwon et al. (2023); Schioppa et al. (2022); Grosse et al. (2023), active learning Chhabra et al. (2024b); Liu et al. (2021b), etc.). Moreover, while some recent works have considered using influence for selecting direct demonstrations (Nguyen and Wong, 2023; Van et al., 2024), neither of them has considered their effect on inductive bias selection in the indirect ICL setting, which is the focus of our work.
@@ 6 Conclusion
We formalize a new paradigm for generalized In-Context Learning, which we term Indirect In-Context Learning. We analyze two different real-world Indirect ICL settings and propose effective demonstration selection strategies for these scenarios. We explore using Influence Functions (IFs) to leverage the informativeness of the samples in the demonstration pool and the models’ task inductive bias. We find that combining a surrogate model-based IF approach with BertScore performs better when there are an overwhelming majority of irrelevant tasks in the candidate pool. We also find that reweighting the surrogate model-based IF scores with traditional metric scores can be helpful in the case of Noisy ICL. Future work will aim to augment the Pretrained Gradient approach by finetuning the LLMs.
@@
Appendix A Integration to Practical Workflows
We provide several use cases of Indirect ICL in real-world scenarios. Here are a few examples:
A.1 Practical Applications of Indirect ICL
  • Enhancing prompt performance at test time: If an LLM service provider needs to use in-context learning (ICL) to improve prompt performance during test time, they may not know the precise task beforehand or during inference (e.g., a novel task requested by a user in real-time). Indirect ICL and our proposed methods can improve performance by selecting relevant demonstrations from a task-agnostic pool of labeled data (i.e., the MoT setting), ensuring the model can adapt to various scenarios even when task-specific labeled samples (direct supervision) are not available.
  • Medical diagnosis: Indirect ICL can be used to diagnose rare medical conditions based on symptoms. Since such conditions are rare, demonstrations for these specific cases are often unavailable. However, the model can learn diagnostic reasoning patterns from more common conditions with overlapping symptoms, improving accuracy for the rare cases.
  • Code generation for obscure programming languages: Indirect ICL can aid in generating code for rarely-used or proprietary programming languages. Demonstrations from code generation tasks in related languages with similar structures can be leveraged, enabling the model to generalize and perform well in these low-resource scenarios.
  • Ideology Estimation from Underrepresented Contexts: We can use our paradigm to estimate political ideology, or any other sort of text classification, from text in an underrepresented cultural or linguistic context. We can use demonstrations from ideology estimation in well-represented contexts such as Western political texts. The can transfer learned associations between linguistic cues and ideological stances, adapting them to the new context.
These examples highlight just a few of the practical applications of indirect ICL, particularly in low-resource settings.
Appendix B DataInf Method of Computing Influence Functions
DataInf is an efficient method for estimating influence functions in deep neural networks, especially in parameter-efficient fine-tuning settings such as LoRA. Traditional influence function computation requires inverting large Hessian matrices, which is impractical for large models. DataInf addresses this by introducing a closed-form approximation that significantly reduces both computational and memory costs.
The key approximation is to swap the order of the matrix inversion and the average calculations as below:
$\left(\frac{1}{n}\sum_{i=1}^{n}\nabla_{\theta_{\ell}}\ell_{i}\nabla_{\theta_{\ell}}\ell_{i}^{\top}+\lambda_{\ell}I_{d_{\ell}}\right)^{-1}\;\approx\;$
$\frac{1}{n}\sum_{i=1}^{n}\left(\nabla_{\theta_{\ell}}\ell_{i}\nabla_{\theta_{\ell}}\ell_{i}^{\top}+\lambda_{\ell}I_{d_{\ell}}\right)^{-1}$
Using the Sherman-Morrison formula, the inverse in each term can be computed analytically, allowing a closed-form estimate of the influence:
$\mathcal{I}_{\text{D}}(x_{k},y_{k})=\sum_{\ell=1}^{L}\frac{1}{\lambda_{\ell}}\left(\frac{1}{n}\sum_{i=1}^{n}\frac{L_{\ell,i}}{\lambda_{\ell}+L_{\ell,ii}}L_{\ell,ik}-L_{\ell,k}\right)$
This formulation enables DataInf to scale to large models by avoiding iterative solvers and full Hessian storage, making it practical for real-world LLM applications.
Appendix C Indirect ICL Results for LLM based Influence
C.1 Full Results for Mixture of Tasks
@@
Following are the full results for the Mixture of Tasks setting. For $k=3$ shots in Tables 3, 4, and 5. For $k=5$ shots in Table 6.
@@
Table 3: Performance across different datasets and demonstration selection methods with $k=3$ shots. The datasets are sampled from sub-tasks of the MMLU and BigBench datasets for Llama-2-13b-chat.
@@
Dataset Rand BSR Cos BM25 PreD SurD Sur[D,BSR] Sur[D,Cos] medical-genetics 80.60 87.00 84.00 79.00 76.00 81.00 86.00 82.00 prof-psychology 68.30 70.00 73.00 65.50 65.00 68.50 72.00 68.50 formal-logic 60.16 59.52 60.32 58.73 61.11 59.52 57.14 55.56 moral-disputes 81.00 78.00 79.50 80.50 78.50 76.00 80.50 76.50 public-relations 72.18 79.09 80.91 73.64 71.82 75.45 79.09 79.09 comp-security 76.80 76.00 80.00 76.00 76.00 80.00 76.00 76.00 astronomy 80.26 80.26 78.95 80.92 74.34 79.61 80.26 78.95 abstract-algebra 57.00 58.00 62.00 55.00 57.00 47.00 72.00 72.00 nutrition 75.50 77.50 79.00 78.00 77.50 77.00 79.50 81.50 high-school-biology 76.70 76.50 78.00 76.50 73.50 79.00 80.50 76.50 formal-fallacies 47.25 52.00 50.00 49.50 47.00 56.50 47.50 50.00 tracking-3 40.20 44.00 39.00 45.00 40.00 37.00 43.50 39.00 Average 68.00 69.82 70.39 68.19 66.48 68.04 71.16 69.63
@@
Table 4: Performance across different datasets and demonstration selection methods with $k=3$ shots. The datasets are sampled from sub-tasks of the MMLU and BigBench datasets for Mistral-7b-v3
@@
Table 5: Performance across different datasets and demonstration selection methods with $k=3$ shots. The datasets are sampled from sub-tasks of the MMLU and BigBench datasets for Zephyr-7b-beta
@@
Table 6: Performance across different datasets and demonstration selection methods with $k=5$ shots for Llama-2-13b-chat
@@
C.2 TracIn Results
Here we provide results for the TracIn method of Influence Computation for $k=3$ shots in Tables 7, 8, and 9. We also provide results for $k=5$ shots in Table 10.
Table 7: Performance across different datasets and different TracIn Influence methods with $k=3$ shots for Llama2-13b-chat
@@ C.2 TracIn Results
Table 8: Performance across different datasets and different TracIn Influence methods with $k=3$ shots for Mistral-7b-v0.3
@@
Table 9: Performance across different datasets and different TracIn Influence methods with $k=3$ shots for Zephyr-7b-beta
@@
Table 10: Performance across different datasets and different TracIn Influence methods with $k=5$ shots for Llama2-13b-chat.
@@
C.3 Pretrained Gradient Results
Here we provide results for pretrained gradients method of computing IF. These can be found in for $k=3$ shots in Tables 11, 12 and 13 and for $k=5$ shots in Table 14.
@@ C.3 Pretrained Gradient Results
Table 12: Performance across different datasets and Pre-training based demonstration selection methods with $k=3$ shots for Mistral-7b-v0.3.
@@
Table 13: Performance across different datasets and Pre-training based demonstration selection methods with $k=3$ shots for Zephyr-7b-beta.
@@
Table 14: Performance across different datasets and Pre-training based demonstration selection methods with $k=5$ shots for Llama2-13b-chat.
@@
C.4 Results using a different surrogate model
We present results where DeBERTa-v3-Large replaces RoBERTa-Large as the surrogate model. Evaluated on Llama2-13B-Chat with $k=3$ shots. We compare the best-performing baseline Sur[D,BSR] with BSR in Table 15.
The results indicate that the DeBERTa surrogate model outperforms BSR. However, it is important to note that, given the inclusion of only three shots in the prompt—where the overwhelming majority of demonstrations are unrelated to the test task—achieving significant improvements remains challenging.
Table 15: Performance comparison between BSR and Sur[D,BSR] with DeBERTa as the surrogate model with ($k=3$ shots) for Llama2-13b-chat.
Dataset BSR Sur[D,BSR] medical-genetics 87.00 85.00 prof-psychology 70.00 72.50 formal-logic 59.52 59.52 moral-disputes 78.00 84.50 public-relations 79.09 76.36 comp-security 76.00 77.00 astronomy 80.26 80.26 abstract-algebra 58.00 59.00 nutrition 77.5 79.00 high-school-biology 76.50 79.00 formal-fallacies 52.00 45.5 tracking-3 44.00 46.50 Average 69.82 70.30
C.5 Results on Qwen2.5-3b
Here we present results on Qwen2.5-3b Team (2024) for $k=3$ shots on the Mixture of Tasks setting. We compare BSR with our two IF based BSR methods Sur[D,BSR] and Pre[D,BSR]. We see in Table 16 that for an advanced model like Qwen2.5-3b, the unfinetuned pretrained gradients slightly outperform the finetuned gradients from the RoBERTa model.
Table 16: Performance comparison between BSR, Sur[D,BSR], and Pre[D,BSR] with ($k=3$ shots) for Qwen2.5-3b.
Dataset BSR Sur[D,BSR] Pre[D,BSR] medical-genetics 88 90 87 prof-psychology 80 82.5 81.5 formal-logic 71.43 66.67 68.25 moral-disputes 83.5 80 82.5 public-relations 79.09 79.09 80.91 comp-security 84 87 85 astronomy 85.53 88.16 88.82 abstract-algebra 64 63 63 nutrition 84.5 85.5 85 high-school-biology 87.5 87 89 formal-fallacies 49 50 50.5 tracking-3 48 48 49 Average 75.38 75.57 75.87
C.6 Ablation analysis on the benefits of combining IF with semantic similarity methods
We conduct an ablation analysis to examine the benefit of using BertScore-Recall and Cosine Similarity in conjunction with our IF-based methods. We want to quantify how effective these methods are at selecting the same training task demonstration as the test task in the 3 shots selected (e.g if the test task is from MMLU-abstract-algebra, then the demonstration retrieved is also from the MMLU-abstract-algebra train set in our demonstration pool). We find that BertScore-Recall selects the same task 33.90% of the time and Cosine Similarity selects the same task 36.03% of the time. In our MoT approach, we initially retrieve the top $2k$ candidate demonstrations using similarity-based methods, and subsequently re-rank and select the final $k$ examples based on IF scores. This staged retrieval process effectively narrows the pool to semantically relevant demonstrations, enabling the IF-based re-ranking to more precisely exploit the model’s inductive biases.
C.7 Qualitative Analysis
For MoT, to understand the merits of our method, we compare demonstrations selected via BSR and Sur[D,BSR] on the MMLU-abstract-algebra dataset:
Example 1:
Q: Compute the product in the given ring. $(2,3)(3,5)$ in $\mathbb{Z}_{5}\times\mathbb{Z}_{9}$
Options: (B) (3,1)  (C) (1,6)
BSR Shots.
  1. 1.
    Q: Statement 1 | Every element of a group generates a cyclic subgroup of the group. Statement 2 | The symmetric group $S_{10}$ has 10 elements. Options: (A) True, True  (C) True, False Answer: (C)
  2. 2.
    Q: Statement 1 | Every function from a finite set onto itself must be one-to-one. Statement 2 | Every subgroup of an abelian group is abelian. Options: (A) True, True  (D) False, True Answer: (A)
  3. 3.
    Q: How many attempts should you make to cannulate a patient before passing the job on to a senior colleague, according to the medical knowledge of 2020? Options: (A) 4  (B) 3  (C) 2  (D) 1 Answer: (C)
Sur[D,BSR].
  1. 1.
    Q: Statement 1 | Every function from a finite set onto itself must be one-to-one. Statement 2 | Every subgroup of an abelian group is abelian. Options: (A) True, True  (D) False, True Answer: (A)
  2. 2.
    Q: Olivia used the rule "Add 11" to create the number pattern shown below: 10, 21, 32, 43, 54. Which statement about the number pattern is true? Options: (B) The number pattern will never have two even numbers next to each other. (D) If the number pattern started with an odd number, then the pattern would have only odd numbers in it. Answer: (B)
  3. 3.
    Q: Tomorrow is 11/12/2019. What is the date one year ago from today in MM/DD/YYYY format? Options: (B) 11/11/2018  (C) 08/25/2018 Answer: (B)
We can see that while BSR selects more semantically relevant samples, Sur[D,BSR]’s selected shots guide the model toward the correct answer (C) instead of (B) by encouraging more structured reasoning.
Example 2:
Q: Statement 1 | If $R$ is an integral domain, then $R[x]$ is an integral domain. Statement 2 | If $R$ is a ring and $f(x)$ and $g(x)$ are in $R[x]$, then
$\deg(f(x)g(x))=\deg f(x)+\deg g(x).$
Options: (C) True, False  (B) False, False
BSR Shots.
  1. 1.
    Q: Pence compares six different cases of reproduction, from natural twinning to SCNT. What conclusion does he draw from this comparison? Options: (A) SCNT is not a different kind of reproduction because there are no morally relevant differences between it and other permissible means of reproduction. (B) Because there is a low risk of harm for natural twinning, there will be a low risk of harm for SCNT. (C) Both A and B (D) Neither A nor B Answer: (A)
  2. 2.
    Q: Statement 1 | Every element of a group generates a cyclic subgroup of the group. Statement 2 | The symmetric group $S_{10}$ has 10 elements. Options: (A) True, True  (C) True, False Answer: (C)
  3. 3.
    Q: Statement 1 | Every function from a finite set onto itself must be one-to-one. Statement 2 | Every subgroup of an abelian group is abelian. Options: (A) True, True  (D) False, True Answer: (A)
Sur[D,BSR].
  1. 1.
    Q: Statement 1 | Every function from a finite set onto itself must be one-to-one. Statement 2 | Every subgroup of an abelian group is abelian. Options: (A) True, True  (D) False, True Answer: (A)
  2. 2.
    Q: Select the best translation into predicate logic. George borrows Hector’s lawnmower. ($g$: George; $h$: Hector; $l$: Hector’s lawnmower; $Bxyx$: $x$ borrows $y$ from $z$). Options: (A) $B_{lgh}$ (B) $B_{hlg}$ (C) $B_{glh}$ (D) $B_{ghl}$ Answer: (C)
  3. 3.
    Q: Statement 1 | Every element of a group generates a cyclic subgroup of the group. Statement 2 | The symmetric group $S_{10}$ has 10 elements. Options: (A) True, True  (C) True, False Answer: (C)
Here again, BSR selects the more semantically relevant shots (with the top three shots ordered in ascending order of relevance), while Sur[D,BSR] selects less semantically similar but more influential shots, which ultimately improves model performance.
Appendix D Extended Noisy ICL Results
D.1 Varying Noise Levels
We also test whether our method can perform well on varying noise levels in the dataset. To test this, we create 2 datasets of MRPC with 10% and 30% noise added. As seen in Table 17, in both the cases IF Averaging outperformed other baselines. With the DataInf configuration performing better for the 10% noise dataset and the LiSSA configuration performing better for the 30% noise dataset.
D.2 Varying Hyperparameters
@@
Here we provide results for different IF pruning and IF averaging hyperparameters that we tested with varying levels of noise in Table 18 and Table 19.
@@
Method MRPC 0.2 QNLI 0.2 SST2 0.2 QQP 0.2 PRU-0.2 Rand[Cos] 68.1 68.6 86.6 70.0 Sur[D,Cos] 67.7 67.2 86.2 70.0 Sur[L,Cos] 71.3 65.8 86.8 67.2 Rand[BSR] 69.1 72.2 85.4 69.8 Sur[D,BSR] 70.3 67.4 81.8 71.6 Sur[L,BSR] 70.1 66.4 85.2 70.8 PRU-0.3 Rand[Cos] 70.1 68.4 87.0 67.4 Sur[D,Cos] 68.8 69.2 84.6 70.6 Sur[L,Cos] 70.6 68.2 86.6 72.6 Rand[BSR] 70.8 65.8 86.0 70.8 Sur[D,BSR] 70.6 67.4 84.8 72.0 Sur[L,BSR] 69.4 68.2 87.4 68.4 AVG-0.4 Sur[D,Cos] 75.7 71.4 91.6 62.6 Sur[L,Cos] 73.3 67.2 90.6 75.2 Sur[D,BSR] 74.3 68.8 89.8 65.2 Sur[L,BSR] 56.6 72.8 78.0 73.0 AVG-0.6 Sur[D,Cos] 73.5 69.2 87.2 66.2 Sur[L,Cos] 72.1 68.6 84.6 70.8 Sur[D,BSR] 73.5 69.2 94.2 71.8 Sur[L,BSR] 72.3 65.8 94.2 75.4
D.3 Effectiveness of IFs in identifying Mislabeled Data
We conduct a toy experiment to evaluate the effectiveness of IF-based methods in detecting noisy samples. We introduce 20% noise to the datasets and compute IF values using the Surrogate Model approach. We then calculate the percentage of noisy samples in the top 100 values selected by our IF methods. Results are presented in Table 20
Table 20: Percentage of noisy samples in the top 100 values selected by IF methods.
Dataset DataInf LiSSA
MRPC 83% 66%
QNLI 54% 86%
QQP 79% 95%
SST-2 90% 96%
As shown in the table, IF-based methods are highly effective in identifying mislabeled data, significantly aiding demonstration selection in Noisy ICL.
Appendix E Memory Consumption
To analyze the added computational costs associated with IFs, we calculate the maximum GPU memory consumption while performing demonstration selection with the Pretrained Gradients-DataInf PreD and Surrogate Model-DataInf SurD methods. The experiments are performed on 4 NVIDIA RTX 6000 Ada Generation GPUs. The maximum memory consumption for Pre-D was 18,188 MiB, while for Sur-D it was 7,998 MiB. These memory requirements are relatively modest, and the use of IFs can be justified given the benefits they provide.
Appendix F Scalability to Large Models
We compare the time it takes to extract test set gradients and compute influence scores. We compute IF via the DataInf method, comparing RoBERTa-Large (125 million parameters), Llama2-13b-chat (13 billion parameters), and Llama2-70b-chat (70 billion parameters) on the MMLU-moral-disputes dataset with 200 test samples.
Table 21: Time taken for extracting test gradients and computing IF across different models.
Model Test Gradients (s) Computing IF (s) RoBERTa 7.447 35.72 Llama-2-13b-chat 68.5 4.69 Llama-2-70b-chat 257.64 8.81
The relationship between model size and inference time grows sublinearly, with time increasing at roughly the square root of the model size. We also see that it takes longer to compute the IF in the RoBERTa model due to the fine-tuning process [^9] We fine-tuned the RoBERTa model and computed IF using gradients from its LoRA-adapted components. In contrast, no fine-tuning was performed on the larger LLaMA2-13B-Chat model; consequently, only LayerNorm weights produced non-zero gradients. This significantly reduced the parameter space and computational cost of IF estimation, as IFs depend on the inverse Hessian with respect to model parameters. .
Additionally, the time required to compute IF using the TracIn method on Llama-2-13b-chat is just $\mathbf{3.576\times 10^{-6}}$ seconds. This highlights the significant speed advantage offered by TracIn.
We would like to emphasize that practitioners have the flexibility to choose between our models and methods based on their specific needs. If computational efficiency is the priority, the significantly faster surrogate model approach can be used. Conversely, if high accuracy is desired and compute is not a concern, a fine-tuned LLM is a better alternative.
Appendix G Scalability to large datasets
For larger datasets, we compare the time taken to extract test gradients and compute IF for 100 samples, 200 samples, and 1000 samples in Table 22.
Table 22: Time taken for computing test gradients and influence functions (IF) across different models and sample sizes.
Model & Samples Test Gradients (s) Computing IF (s) RoBERTa (100 Samples) 6.7 26.4 RoBERTa (200 Samples) 7.4 35.7 RoBERTa (1000 Samples) 67.8 273.7 Llama-2-13b-chat (100 Samples) 41.0 3.6 Llama-2-13b-chat (200 Samples) 68.5 4.7 Llama-2-13b-chat (1000 Samples) 410.8 35.4
A 10x increase in sample size corresponds to an approximately 10x increase in computational time, indicating a linear relationship between sample size and computational time.
[…]
Appendix H IF-based demonstrations as a backdoor defense strategy to mitigate backdoor attacks
To evaluate the robustness of our influence-based demonstration selection method under a different class of adversarial noise, we extend it to a task-agnostic backdoor defense setting. This scenario reflects practical constraints where task-specific labeled data may be unavailable at inference time, yet a broader pool of related examples is accessible.
We use our Sur[D,BSR] strategy to the SST-2 dataset Wang (2018), which has been poisoned with two distinct forms of backdoor attacks: (1) AddSent Dai et al. (2019), where a semantic trigger is introduced by inserting an innocuous sentence (e.g., “I watched this 3D movie last weekend”), and (2) Style Qi et al. (2021), a style-based attack that performs backdoor poisoning through text style transfer (e.g., transforming text into a Biblical style).
We evaluate the Attack Success Rate (ASR), defined as the percentage of non-target label test instances that are misclassified as the target label when evaluated on a poisoned dataset. For comparison, we consider a task-aware demonstration selection baseline Diao et al. (2023), which selects 5 demonstrations from a clean, task-specific dataset. All experiments are conducted using the LLaMA3-8B model Grattafiori et al. (2024). We present our results in Table 23.
SST-2 AddSent Style
ASR
No Defense 100.00 98.68
Task-Aware 69.41 32.02
Sur[D,BSR] 58.22 10.86
Table 23: Defense results on AddSent and Style backdoor attacks. ASR: Attack Success Rate (the lower the better).
Our results show that our IF-based indirect-ICL paradigm can effectively mitigate various types of backdoor attacks. This reveals that even without task-specific data, our results demonstrate that the IF-based indirect-ICL paradigm effectively mitigates a range of backdoor attacks, showing a 32.89% reduction in ASR, on average. Notably, even in the absence of task-specific data, demonstrations selected based on a model’s inductive bias can provide a strong defense mechanism, highlighting the potential of task-agnostic strategies as an effective defense mechanism against backdoor attacks.